Goodmorning, I need help with this
Solve the system by graphing, addition and substitution methods.
3x + y = –1
9x + 3y = –3
Thank you

Answer by rapaljer(4667)

3x + y = –1
9x + 3y = –3


Multiply both sides of the first equation by -3 to eliminate the y terms.
-3*(3x + y) = (-3)*–1
9x + 3y = –3

-9x -3 y = 3
9x + 3y = –3

When you add the equations together you get 0=0, which is ALWAYS TRUE. Any time you get 0=0, it means that the equations actually represent the SAME LINE. The entire line is the solution.

R^2